If you obtain a $200,000. 30 year fixed rate mortgage at 6% with two points at closing but sold your house and paid off the mortgage after five years. what were the effective interest rate be over this five-year period ?
Answer #1

SEE THE IMAGE. ANY DOUBTS, FEEL FREE TO ASK. THUMBS UP PLEASE

Home nert Page Layout Formulas Data Review View dd-Ins Cut Σ AutoSum ー E ゴWrap Text ta copy. в 1 프· ー· 鱼, Δ. :r-ㄧ 逻锂函Merge & Center. $, % , 弼,8 Conditional Format eCell Insert Delete Formsat Paste Sort &Find & 2 ClearFe Select Edting Format Painter Formatting, as Table w styles. ▼ ㆆ ▼ Clipboard Font Alignment Number Cells Y273 AB AC AD AE AF 273 274 275 276 277 278 279 280 281 282 283 284 285 286 287 288 289 290 291 LOAN AMOUNT 2 DISCOUNT POINTS RATE LIFE PERIODS IN A YEAR PMT = 200000 4000 (200000X 296) 6.00% 30 12 1,199.10 EXCEL-PMT(AA276/AA278,AA277*AA278,-AA274) LOAN BALANCE AFTER S 1,86,108.71 EXCEL-PV(AA276/AA278,25 *AA278,-AA279) WITH DISCOUNT POINTS TOTAL LOAN AMOUNT $1,90,108.71 (186108.714000) MONTHLY COST 0.5293% EXCEL RATE(5 AA278,-AA279,AA274,-AA286) 6.54% (1 +0.5293%)^12-1 EAR = (1+ r/m)^m-1 EFFECTIVE ANNUAL COST r/m = MONTHLY RATE, m-12 CALCULATOR | LOAN OPTIONS fv. ann CAP STRU VALUE LOCK BOX DILUTION DOLLAR COST AVG MORTGAGE EXPO Sheeti Sheet2 Shel 福 130% 04:21 31-01-2019